hi i was wondering if you can give me an exsample of finding the midpoint of 2 joining points my equation is E (3,5) F(8,9). thank you
Answer by Alan3354(69443)

That's not an equation. It's 2 points.
Equations have an equal sign. equal - equate - equation
-------------------------------------------------------
To find the midpoint, find the average of x and y separately.
x: (3+8)/2 = 11/2 or 5.5
y: (5+9)/2 = 7
MP at (5.5,7)